OpenStudy (anonymous):

8! - 6! ------ 3! x 5!

OpenStudy (anonymous):

8!/3!*5! - 6!/3!*5! 8*7*6/6 - 6/6 56 - 1 55

OpenStudy (anonymous):

couldn't understand... kindly explain a bit

OpenStudy (unklerhaukus):

\[\frac{8!-6!}{3!\times5!}\]\[=\frac{8!}{3!\times5!}-\frac{6!}{3!\times5}\]\[=\frac{8\times7\times6\times5!}{3!\times5!}-\frac{6\times5!}{3!\times5!}\]\[=\frac{8\times7\times6}{3!}-\frac{6}{3!}\]

OpenStudy (unklerhaukus):

\[=\frac{8\times7\times6}{3\times2\times1}-\frac{6}{3\times2\times1}\]\[=\frac{8\times7\times6}{6}-\frac{6}{6}\]\[=8\times7-1\]

OpenStudy (anonymous):

Thanx... got it